OpenStudy (anonymous):

Write the equation of a circle with a center at (–7, –7) and a radius of 7. (x + 7)2 + (y + 7)2 = 14 (x + 7)2 + (y + 7)2 = 49 (x – 7)2 + (y – 7)2 = 49 (x + 7)2 + (y + 7)2 = 7

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Sorry I wasn't on :( But the equation of a circle is \[\large (x - h)^2 + (y - k)^2= r^2\] where \(\large h,k\) is the center of a circle and r is the radius...so here \[\large (x + 7)^2 + (y + 7)^2 = 7^2\] \[\large (x + 7)^2 + (y + 7)^2 = 49\]

OpenStudy (anonymous):

no its okay :) thankn you so much @johnweldon1993 The equation x equals one-twenty-fourth times y squared represents a __________. circle ellipse parabola hyperbola

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

Circles M and K are congruent, QR is congruent to LN and OP is congruent to VW. Find y. http://curriculum.kcdistancelearning.com/courses/ALG2x-HS-A09/b/assessments/Semester2Exam/A2_Semester_B_Final_102q.gif 3 4 -4 -3

OpenStudy (anonymous):

What is the r value of the following geometric sequence? -81, 54, -36, 24,… negative five-ninths negative two-thirds negative five-sixths negative three-fourths

OpenStudy (anonymous):

What is the probability of picking an “o” in the word “footloose”? 4/5 1/9 1/5 4/9

OpenStudy (anonymous):

What is the equation of the ellipse with foci (0, 6), (0, -6) and co-vertices (2, 0), (-2, 0)? x squared divided by 1 plus y squared divided by 36 equals 1 x squared divided by 1 minus y squared divided by 36 equals 1 x squared divided by 1 plus y squared divided by 40 equals 1 x squared divided by 1 minus y squared divided by 40 equals 1